# customstyle linecolor=gray!100, linewidth=3pt, innerleftmargin=3pt, topline=false, rightline=false, bottomline=false, leftline=true, innerrightmargin=3pt, innertopmargin=3pt, innerbottommargin=3pt, backgroundcolor=gray!15

### v1

@@ 4.3. Metrics
We use two evaluation metrics: $Accuracy^{c}$ proposed by Utango (li2022utango, ) and $Accuracy^{a}$ proposed by Flexeme (parțachi2020flexeme, ). $Accuracy^{c}$ is defined as the percentage of the changed statements that are labeled with a correct cluster/concern in all the statements in a commit:
@@ 4.4. Implementation Details
We utilize the widely adopted and publicly available DeepSeek-V3 model (liu2024deepseek, ) from the Azure OpenAI service [^1] https://learn.microsoft.com/en-us/azure/ai-services/openai/ as the reasoning engine in our experiments, which is also used for the LLM-based baselines. All experiments are conducted under a zero-shot setting. From an initial evaluation on 50 randomly selected tangled commits, we observed an average of 1.27 consultation rounds to reach consensus. Thus, we set the maximum number of rounds $t$ to 3, aligning with ColaCare (wang2025colacare, ) (a medical multi-agent collaborative consultation model), and balancing iterative refinement with computational efficiency. In terms of computational cost, our approach incurs approximately \$2.92 per 100 examples (around ¢2.92 per example), with an average inference time of 125.13 seconds per example [^2] We acknowledge that our approach requires longer inference time compared to prior methods. However, it eliminates the need for manual heuristic rule design, feature engineering, and model training, making it relatively cost-effective given its significant improvement in untangling accuracy..
@@ 5. Experimental Results
Table 1. Performance on the C# Dataset ($Accuracy^{c}$%/$Accuracy^{a}$%)
@@
    ColaUntangleno_comments refers to run our method on the dataset excluding comments.
@@
$LLM_{ZeroShot}$ 66 68 66 50 54 60 51 73 66 57 63
$LLM_{ZeroShotCoT}$ 66 70 68 68 62 62 49 70 65 57 66
ColaUntangleno_comments 70 80 69 77 60 75 67 71 79 71 76
ColaUntangle 70 79 70 78 61 75 66 73 80 72 76
  • Since previous studies do not report $Accuracy^{a}$on the Java dataset, we follow their practice and report only $Accuracy^{c}$to ensure a fair comparison.
@@ 5.1. RQ1: Overall Performance
Table 1 compares metrics on the C# dataset (including nine GitHub projects: CL: Commandline, CM: CommonMark, HF: Hangfire, HU: Humanizer, LE: Lean, NA: Nancy, NJ: Newtonsoft.Json, NI: Ninject, RS: RestSharp). Compared to previous untangling approaches, ColaUntangle outperforms SOTA, improving Barnett et al., Herzig et al., $\delta$-PDG+CV, Flexeme, Utango and HD-GNN by 928.57%, 157.14%, 105.71%, 111.76%, 56.52% and 44.00% for $Accuracy^{c}$; 615.38%, 40.91%, 10.71%, 12.05%, 3.33% and 3.33% for $Accuracy^{a}$. For LLM-Based baselines, ColaUntangle improves $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% on $Accuracy^{c}$, and by 3.33% on $Accuracy^{a}$ for both. Compared to ColaUntangleno_comments, which excludes comment statements, ColaUntangle achieves a 2.86% higher $Accuracy^{c}$, while $Accuracy^{a}$ remains the same.
Table 2 shows the Java dataset comparison (including ten GitHub projects: SB: spring-boot, ES: elasticsearch, RJ: RxJava, GU: guava, RE: retrofit, DU: dubbo, GH: ghidra, ZX: zxing, DR: druid, EB: EventBus), where ColaUntangle improves Barnett et al., Herzig et al., Base-1, Base-2, SmartCommit Utango and HD-GNN by 347.06%, 230.43%, 153.33%, 347.06%, 216.67%, 123.53% and 100.00% for $Accuracy^{c}$. For LLM-based baselines, ColaUntangle improves over $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% respectively, and is the same as ColaUntangleno_comments for $Accuracy^{c}$.
The results show that ColaUntangle consistently outperforms all baseline models, with notable improvements in the $Accuracy^{c}$ metric. Importantly, unlike most existing methods, ColaUntangle can untangle comment statements, which are common cosmetic edits in real-world commits, making it more practical for software development. ColaUntangleno_comments also surpasses all prior approaches, indicating that while comments can slightly enhance performance, our framework remains effective even without them.
Regarding collaborative consultation rounds, the average number of rounds is 1.38: 75.50% of cases converge within one round, 11.53% within two rounds, and 13.16% reach the maximum of three rounds. Notably, in 48.91% of evaluated commits, the explicit and implicit worker agents initially produced non-consensual results, underscoring the importance of collaborative consultation for resolving divergent untangling perspectives.
@@
Figure 6. Average $Accuracy^{c}$ and Average Number of Changed Statements in Tangled Commits of Each Repository
@@
{mdframed}
[style=customstyle] Answer to RQ1: ColaUntangle demonstrates improvements over prior baselines on both datasets, improving $Accuracy^{c}$ by 44% on the C# dataset and 100% on the Java dataset over the best existing baseline. The results highlights its potential for practical application in untangling commits.
@@ 5.2. RQ2: Ablation Study
We conduct an ablation study to evaluate the contribution of each component within ColaUntangle. As shown in Table 3, we define four ablation settings: (1) w/o Info Tools: Removes explicit and implicit contexts extracted from the $\delta$-PDG. (2) w/o Collaborative Consultation: Uses a single LLM model to mine both explicit and implicit dependencies without the multi-agent consultation process. (3) w/o Explicit Worker Agent (W1) and Collaborative Consultation: Retains only the implicit worker agent to perform untangling. Since there is only a single agent remaining, the collaborative consultation process becomes unnecessary, and the reviewer agent is thus excluded. (4) w/o Implicit Worker Agent (W2) and Collaborative Consultation: The converse of (3), retaining only the explicit worker agent.
From the results, we observe that excluding Collaborative Consultation (setting 2) leads to the most significant performance degradation (17.57%) [^3] For the sake of brevity, all mentions of performance degradation in this section refer specifically to a drop in $Accuracy^{c}$, as it better reflects the model’s performance in real-world tasks compared to $Accuracy^{a}$., highlighting its critical role in achieving optimal untangling results. Removing either the explicit or implicit worker agent (settings 3 and 4) also results in significant performance drops (20.27% and 18.92%, respectively). However, it is important to note that these two settings inherently exclude Collaborative Consultation due to the presence of only a single agent, making their performance degradation partly attributable to the absence of collaborative consultation. When we isolate the effect by excluding the performance drop caused by the lack of collaborative consultation, settings 3 and 4 show additional reductions of only 2.70% and 1.35%, respectively. This indicates that while incorporating both explicit and implicit worker agents enhances performance, the collaborative consultation mechanism itself contributes the most substantial improvement in our framework. Removing Info Tools (setting 1) results in a performance decrease of 9.46%, suggesting that the structured context improves our model’s untangling effectiveness. However, compared to the substantial improvements brought by collaborative consultation, Info Tools mainly serves as foundational support by providing essential contextual information rather than driving the core reasoning and decision-making process.
{mdframed}
[style=customstyle] Answer to RQ2: The ablation results suggest that while structured context information and the integration of both explicit and implicit worker agents supports untangling accuracy, the collaborative consultation mechanism play a more critical role in achieving the effectiveness of ColaUntangle.
Table 3. Ablation Study Results for Each Module ($Accuracy^{c}$%/$Accuracy^{a}$%)
@@
w/o Info Tools 64/91 70/92 67/92
w/o Collaboraive 60/90 63/92 61/91
w/o W1 w/o Collaborative 57/89 63/91 59/90
w/o W2 w/o Collaborative 58/90 63/92 60/91
ColaUntangle 72/93 76/94 74/94
@@ 5.3. RQ3: Sensitivity to Different LLMs
To evaluate the generalization capability of our framework, we include both closed-source proprietary models (GPT-4o, GPT-4o-mini from OpenAI, Claude-4-sonnet from Anthropic) and open-source frontier models (DeepSeek-V3 and Qwen3-235b) in our comparison. We compare ColaUntangle instantiated with these five models respectively. The performance presented in Table 4 shows that all these LLMs can effectively untangle commits, with DeepSeek-V3, Claude-4-sonnet and GPT-4o-mini performing slightly better. This suggests that while ColaUntangle is generally robust to the underlying LLM choice, stronger reasoning models still yield marginal performance gains. Furthermore, we analyze the average number of collaborative rounds required for consensus: DeepSeek-V3 (1.38), GPT-4o-mini (1.65), GPT-4o (1.35), Qwen3-235b (1.39), and Claude-4-sonnet (1.77). Notably, DeepSeek-V3 not only attains leading accuracy but also requires fewer consensus rounds compared to most competing models (except GPT-4o), highlighting its efficiency in balancing performance and computational cost.
{mdframed}
[style=customstyle] Answer to RQ3: Our evaluation demonstrates that while all tested LLMs effectively handle commit untangling, DeepSeek-V3 achieves the optimal balance of accuracy and efficiency (fewer consensus rounds).
Table 4. Performance of Different LLMs ($Accuracy^{c}$%/$Accuracy^{a}$%)
@@
GPT-4o-mini 72/93 71/93 72/93
GPT-4o 68/92 74/94 71/93
Qwen3-235b 71/93 73/93 72/93
Claude-4-sonnet 70/93 76/94 73/93
ColaUntangle (DeepSeek-V3) 72/93 76/94 74/94
6. Discussion of Error Cases
In the RQ1 experiments, although ColaUntangle performed well overall, it still misclassified some cases. To support future improvements, we conducted a manual analysis following (fan2024HD-GNN, ), randomly sampling 50 misclassified cases. As shown in Table 5, we grouped the errors into two main types: fine and coarse grouping granularity. Fine grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at finer granularity than the standard answer. In contrast, coarse grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at coarser granularity than the standard answer. Further analysis identified five sub-types that specify the scenarios of these errors.
@@ 6. Discussion of Error Cases
For fine grouping granularity, we observed three representative sub-types: judgment errors (14 cases), composite commit (17 cases), and file formatting (3 cases). Judgment errors (14 cases) occur when ColaUntangle misinterprets the relationship between changes. For example, in Figure 7 A, changes across two files belong to a single concern related to signature revision, but ColaUntangle misclassified them into two concerns by treating lines 9–10 as an isolated function call modification (configure) unrelated to the other file. Similarly, in Figure 7 B, changes at lines 3–5 and 8–9 belong to the same concern (add more logging to jgroups disco), but ColaUntangle separated them because it focused on different logger methods (debug vs. warn). Composite commit (17 cases) occur when the standard atom commit in the dataset actually contains multiple unrelated activities. For instance, Figure 7 C shows a commit that modifies comments for the create function (lines 3–5) and changes the signature of canAcquire, with the two code changes separated by over 240 lines and lacking any logical connection. ColaUntangle classified them into separate concerns, diverging from the dataset’s single-concern annotation. File formatting (3 cases) involve changes unrelated to human edits, such as automatic formatting. As shown in Figure 7 D, the first change in this commit is related to variable revise (not showed in Figure 7 D), the second change adds a newline at the end of the file, which is required by POSIX standards but inconsistently enforced by code editors (posixlinedef, ). Although such changes have no functional impact and are often ignored by developers, ColaUntangle classified them as a separate formatting concern, resulting in misclassification.

### v2

@@ 4.3. Metrics
We use two evaluation metrics: $Accuracy^{c}$ proposed by Utango (li2022utango) and $Accuracy^{a}$ proposed by Flexeme (parțachi2020flexeme). $Accuracy^{c}$ is defined as the percentage of the changed statements that are labeled with a correct cluster/concern in all the statements in a commit:
@@ 4.4. Implementation Details
We construct the PDGs of C# and Java dataset with the assistance of the Flexeme extractor [^1] https://github.com/PPPI/Flexeme and Progex extractor [^2] https://github.com/ghaffarian/progex respectively. We utilize the widely adopted and publicly available DeepSeek-V3 model (liu2024deepseek) from the Azure OpenAI service [^3] https://learn.microsoft.com/en-us/azure/ai-services/openai/ as the reasoning engine in our experiments, which is also used for the LLM-based baselines. All experiments are conducted under a zero-shot setting. All LLMs used in our experiments are accessed via their online APIs, all decoding parameters are kept at their default values. From an initial evaluation on 50 randomly selected tangled commits, we observed an average of 1.27 consultation rounds to reach consensus. Thus, we set the maximum number of rounds $t$ to 3, aligning with ColaCare (wang2025colacare) (a medical multi-agent collaborative consultation model), and balancing iterative refinement with computational efficiency. In terms of computational cost, our approach incurs approximately \$2.92 per 100 examples (around ¢2.92 per example), with an average untangling time of 71.63 seconds per example (Context Building: 9.79 s; Generation of Initial Results: 12.23 s; Synthesis of Preliminary Untangling Results: 14.68 s; Collaborative Consultation Process: 34.93 s) [^4] We acknowledge that our approach requires longer inference time compared to prior methods. However, it eliminates the need for manual heuristic rule design, feature engineering, and model training, making it relatively cost-effective given its significant improvement in untangling accuracy.. We conduct experiments on a Linux server running Ubuntu 22.04 system, equipped with an Intel® Core™ i7 processor 14700k, 32GB RAM.
@@ 5. Experimental Results
Table 1. Performance on the C# Dataset ($Accuracy^{c}$$Accuracy^{a}$%)
@@
    ColaUntangleno_comments refers to running our method on the dataset excluding comments.
@@
$LLM_{ZeroShot}$ 59 66 57 47 47 57 51 65 64 55 59
$LLM_{ZeroShotCoT}$ 61 69 60 66 52 60 48 63 64 56 63
ColaUntangleno_comments 68 66 69 67 60 65 49 71 73 62 67
ColaUntangle 66 72 67 70 53 69 54 72 73 63 69
  • Since previous studies do not report $Accuracy^{a}$% on the Java dataset, we follow their practice and report only $Accuracy^{c}$% to ensure a fair comparison.
@@ 5.1. RQ1: Overall Performance
Table 1 compares metrics on the C# dataset (including nine GitHub projects: CL: Commandline, CM: CommonMark, HF: Hangfire, HU: Humanizer, LE: Lean, NA: Nancy, NJ: Newtonsoft.Json, NI: Ninject, RS: RestSharp). Compared to previous untangling approaches, ColaUntangle outperforms SOTA, improving Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), $\delta$-PDG+CV (parțachi2020flexeme), Flexeme (parțachi2020flexeme), Utango (li2022utango) and HD-GNN (fan2024HD-GNN) by 928.57%, 157.14%, 105.71%, 111.76%, 56.52% and 44.00% for $Accuracy^{c}$; 615.38%, 40.91%, 10.71%, 12.05%, 3.33% and 3.33% for $Accuracy^{a}$. For LLM-Based baselines, ColaUntangle improves $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.67% and 16.39% on $Accuracy^{c}$, and by 3.33% on $Accuracy^{a}$ for both. Compared to ColaUntangleno_comments, which excludes comment statements, ColaUntangle achieves a 2.86% higher $Accuracy^{c}$, while $Accuracy^{a}$ remains the same.
Table 2 shows the Java dataset comparison (including ten GitHub projects: SB: spring-boot, ES: elasticsearch, RJ: RxJava, GU: guava, RE: retrofit, DU: dubbo, GH: ghidra, ZX: zxing, DR: druid, EB: EventBus), where ColaUntangle improves Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), Base-1, Base-2, SmartCommit (shen2021smartcommit) Utango (li2022utango) and HD-GNN (fan2024HD-GNN) by 305.88%, 200.00%, 305.88%, 187.50%, 130.00%, 102.94% and 81.58% for $Accuracy^{c}$. For LLM-based baselines, ColaUntangle improves over $LLM_{ZeroShot}$ and $LLM_{ZeroShotCoT}$ by 16.95% and 9.52% respectively, and improves over ColaUntangleno_comments by 2.99% for $Accuracy^{c}$.
The results show that ColaUntangle consistently outperforms all baseline models, with notable improvements in the $Accuracy^{c}$ metric. Importantly, unlike most existing methods, ColaUntangle can untangle comment statements, which are common cosmetic edits in real-world commits, making it more practical for software development. ColaUntangleno_comments also surpasses all prior approaches, indicating that while comments can slightly enhance performance, our framework remains effective even without them. To be note, the single LLM’s overall accuracy outperforms commit untangling specific baselines, indicating the effectiveness of LLM in capturing the semantic information of tangled commit and executing the untangling task.
Regarding collaborative consultation rounds, the average number of rounds is 1.39: 75.44% of cases converge within one round, 12.11% within two rounds, and 13.55% reach the maximum of three rounds. Notably, in 48.91% of evaluated commits, the explicit and implicit worker agents initially produced non-consensual results, underscoring the importance of collaborative consultation for resolving divergent untangling perspectives.
@@
Figure 5. Average $Accuracy^{c}$ and Number of Changed Statements in Tangled Commits of Each Repository
@@
Answer to RQ1: ColaUntangle demonstrates improvements over prior baselines on both datasets, improving $Accuracy^{c}$ by 44% on the C# dataset and 82% on the Java dataset over the best existing baseline. The results highlights its potential for practical application in untangling commits.
@@ 5.2. RQ2: Ablation Study
We conduct an ablation study to evaluate the contribution of each component within ColaUntangle. As shown in Table 3, we define four ablation settings: (1) w/o Info Tools: Removes explicit and implicit contexts extracted from the PDGs. (2) w/o Collaborative Consultation: Uses a single LLM model to mine both explicit and implicit dependencies without the multi-agent consultation process. (3) w/o Explicit Worker Agent (W1) and Collaborative Consultation: Retains only the implicit worker agent to perform untangling. Since there is only a single agent remaining, the collaborative consultation process becomes unnecessary, and the reviewer agent is thus excluded. (4) w/o Implicit Worker Agent (W2) and Collaborative Consultation: The converse of (3), retaining only the explicit worker agent.
From the results, we observe that excluding Collaborative Consultation (setting 2) leads to the most significant performance degradation (16.90%) [^5] For the sake of brevity, all mentions of performance degradation in this section refer specifically to a drop in $Accuracy^{c}$, as it better reflects the model’s performance in real-world tasks compared to $Accuracy^{a}$., highlighting its critical role in achieving optimal untangling results. Removing either the explicit or implicit worker agent (settings 3 and 4) also results in significant performance drops (18.99% and 19.72%, respectively). However, it is important to note that these two settings inherently exclude Collaborative Consultation due to the presence of only a single agent, making their performance degradation partly attributable to the absence of collaborative consultation. When we isolate the effect by excluding the performance drop caused by the lack of collaborative consultation, settings 3 and 4 show additional reductions of only 2.09% and 2.82%, respectively. This indicates that while incorporating both explicit and implicit worker agents enhances performance, the collaborative consultation mechanism itself contributes the most substantial improvement in our framework. Removing Info Tools (setting 1) results in a performance decrease of 8.45%, suggesting that the structured context improves our model’s untangling effectiveness. However, compared to the substantial improvements brought by collaborative consultation, Info Tools mainly serves as foundational support by providing essential contextual information rather than driving the core reasoning and decision-making process.
Answer to RQ2: The ablation results suggest that while structured context information and the integration of both explicit and implicit worker agents supports untangling accuracy, the collaborative consultation mechanism play a more critical role in achieving the effectiveness of ColaUntangle.
Table 3. Ablation Study Results for Each Module ($Accuracy^{c}$$Accuracy^{a}$%)
@@
w/o Info Tools 64/91 65/91 65/91
w/o Collaborative 60/90 59/91 59/91
w/o W1 w/o Collaborative 57/90 54/90 56/90
w/o W2 w/o Collaborative 58/90 55/92 57/91
ColaUntangle 72/93 69/93 71/93
@@ 5.3. RQ3: Sensitivity to Different LLMs
To evaluate the generalization capability of our framework, we include both closed-source proprietary models (GPT-4o, GPT-4o-mini from OpenAI, Claude-4-sonnet from Anthropic) and open-source frontier models (DeepSeek-V3 and Qwen3-235b) in our comparison. We compare ColaUntangle instantiated with these five models respectively. The performance presented in Table 4 shows that all these LLMs can effectively untangle commits, with DeepSeek-V3 performing slightly better. This suggests that while ColaUntangle is generally robust to the underlying LLM choice, stronger reasoning models still yield marginal performance gains. Furthermore, we analyze the average number of collaborative rounds required for consensus: DeepSeek-V3 (1.39), GPT-4o-mini (1.65), GPT-4o (1.34), Qwen3-235b (1.41), and Claude-4-sonnet (1.73). Notably, DeepSeek-V3 not only attains leading accuracy but also requires fewer consensus rounds compared to most competing models (except GPT-4o), highlighting its efficiency in balancing performance and computational cost.
Answer to RQ3: Our evaluation demonstrates that while all tested LLMs effectively handle commit untangling, DeepSeek-V3 achieves the optimal balance of accuracy and efficiency (fewer consensus rounds).
Table 4. Performance of Different LLMs ($Accuracy^{c}$$Accuracy^{a}$%)
@@
GPT-4o-mini 72/93 64/91 69/93
GPT-4o 68/92 67/93 67/93
Qwen3-235b 71/93 66/92 69/92
Claude-4-sonnet 70/93 68/92 69/93
ColaUntangle (DeepSeek-V3) 72/93 69/93 71/93
6. Discussion
Although ColaUntangle performs well overall, examining both misclassified and correctly untangled commits provides a deeper understanding of its strengths and limitations, informing potential directions for future work.
6.1. Error Cases
In the RQ1 experiments, although ColaUntangle performed well overall, it still misclassified some cases. To support future improvements, we conducted a manual analysis following Fan et al. (fan2024HD-GNN), randomly sampling 50 misclassified cases. As shown in Table 5, we grouped the errors into two main types: fine and coarse grouping granularity. Fine grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at finer granularity than the standard answer. In contrast, coarse grouping granularity refers to ColaUntangle untangles a tangled commit into concerns at coarser granularity than the standard answer. Further analysis identified five sub-types that specify the scenarios of these errors.
@@ 6.1. Error Cases
For fine grouping granularity, we observed three representative sub-types: judgment errors (14 cases), composite commit (17 cases), and file formatting (3 cases). Judgment errors (14 cases) occur when ColaUntangle misinterprets the relationship between changes. For example, in Figure 6 A, changes across two files belong to a single concern related to signature revision, but ColaUntangle misclassified them into two concerns by treating lines 9–10 as an isolated function call modification (configure) unrelated to the other file. Similarly, in Figure 6 B, changes at lines 3–5 and 8–9 belong to the same concern (add more logging to jgroups disco), but ColaUntangle separated them because it focused on different logger methods (debug vs. warn). Composite commit (17 cases) occur when the standard atom commit in the dataset actually contains multiple unrelated activities. For instance, Figure 6 C shows a commit that modifies comments for the create function (lines 3–5) and changes the signature of canAcquire, with the two code changes separated by over 240 lines and lacking any logical connection. ColaUntangle classified them into separate concerns, diverging from the dataset’s single-concern annotation. File formatting (3 cases) involve changes unrelated to human edits, such as automatic formatting. As shown in Figure 6 D, the first change in this commit is related to variable revise (not showed in Figure 6 D), the second change adds a newline at the end of the file, which is required by POSIX standards but inconsistently enforced by code editors (posixlinedef). Although such changes have no functional impact and are often ignored by developers, ColaUntangle classified them as a separate formatting concern, resulting in misclassification.
@@
6.2. Correct Cases
To further validate the effectiveness of our approach, we randomly selected a representative case from the dataset and conducted a comparative evaluation of ColaUntangle, UTango (li2022utango), and a single LLM. We selected UTango as a key comparator because it partially shares common structural data sources (utilizing $\delta$-PDG, which informs our structured code change) and its methodology touches upon the concepts of explicit and implicit factors. Specifically, UTango employs a sequential, two-stage process: extracting feature embeddings from the $\delta$-PDG for initial clustering (explicit context), followed by a code-clone detection process to refine groupings based on similarity (implicit dependency). This sequential approach is a critical point of contrast with our collaborative mechanism. We selected the single LLM (DeepSeek-V3) as a second comparator to isolate and intuitively observe the performance of LLM on direct reasoning. This baseline operates without access to the structured explicit/implicit contexts or the rule-based guidance provided to our multi-agent framework. Through this comparative setting, we aim to highlight the contribution of ColaUntangle’s multi-agent reasoning and context-aware mechanisms to improving untangling accuracy and interpretability.
Figure 8. Untangling Result of Different Model
As shown in Figure 8, the commit contained three distinct changes: a file update, and the addition of two new functions, SetJobParameter and CreateJob. While UTango and the single LLM failed to untangle this commit correctly, our framework successfully generated the correct answer, which is not a trivial task due to the subtle dependencies between the code changes. UTango failed for two key reasons: it could not identify the first code change because the $\delta$-PDG representation was incomplete (lacking the import statement at the file’s beginning), and it incorrectly grouped the two new functions together based on a structural similarity detected by its code clone detection process, ignoring their functional independence. Similarly, the single LLM also grouped the two functions together, explaining its decision based on a generic principle of separating “cosmetic edits from functional changes,” but it failed to discern the distinct purposes of the two new functions.
Figure 9. Untangling Process of ColaUntangle
The effectiveness of ColaUntangle lies in its underlying process, as shown in our step-by-step analysis (Figure 9). The Explicit Agent, using the explicit contexts and original commit diff, correctly identified the two new functions as separate entities, while the Implicit Agent’s semantic analysis noted their structural and semantic similarities. The Reviewer Agent then synthesized and balanced these two viewpoints, ultimately generating a correct grouping that preserved the functional independence of the new functions. The collaborative consultation process then led all agents to agree on this correct grouping. This example illustrates the combined power of our key ideas: the LLM’s superior semantic understanding, its ability to deal with cosmetic edits and other non-functional changes, and its capacity to use the explicit and implicit contexts as an active component to infer subtle dependencies. The collaborative framework ensures a balanced and accurate result by reconciling the distinct explicit and implicit viewpoints, which together make ColaUntangle’s performance superior to that of prior work and single-model baselines.